OpenStudy (anonymous):

Find the product of (x + 7)2. x2 − 14x + 49 x2 + 14x + 49 x2 − 49 x2 + 49

jimthompson5910 (jim_thompson5910):

Hint: use the formula (A+B)^2 = A^2+2AB+B^2

OpenStudy (anonymous):

im trying to get awnsers not ways how to do it.

jimthompson5910 (jim_thompson5910):

Well only trying to help, not just give you the answers every time. That's not the best way to learn.

OpenStudy (anonymous):

i just got a few questions to do then i can have a good grade and not fail my online school please :/

jimthompson5910 (jim_thompson5910):

alright, why not try working it out and I'll help you along the way?

jimthompson5910 (jim_thompson5910):

first off, what does (x+7)^2 mean?

OpenStudy (anonymous):

because i cant even do the very easyist of alg ... lol umm idk o.o

jimthompson5910 (jim_thompson5910):

Anything squared is equal to that thing multiplied by itself, so x^2 = x*x and (x+7)^2 = (x+7)(x+7)

jimthompson5910 (jim_thompson5910):

Now we FOIL out (x+7)(x+7) (x+7)(x+7) x*x + x*7 + 7*x + 7*7 x^2+7x+7x + 49 x^2+14x+49 So (x+7)(x+7) = x^2+14x+49 Practice with a few to get a better grasp on how FOILing works.